# Unlocking Unlabeled Data: Ensemble Learning with the Hui-Walter Paradigm for Performance Estimation in Online and Static Settings

### v1

@@ 3.2.3 Summary of Experimental Results
Our experiments show that the model’s false positive and false negative rates and prior can be estimated on unknown data sets with the Hui-Walter method. Although we assumed the default distribution $Beta(1,1)$ for the errors and priors, the results from the two experiments are generally promising for rough order of magnitude estimates. With more iterations or a more specific distributional assumption, the results would improve. Results would also likely be more favorable with better trained models as well.
3.3 Hui-Walter Online
@@
Our main goal in this research is to answer how we estimate false positive and false negative rates of binary categorizers to a stream of unlabeled data without available ground truth. Current online machine learning theory still assumes that labels are available for the data in question (Gomes et al., 2019). Figure 6 shows an experiment where we streamed the Wisconsin Breast Cancer Data set with the river python library for online machine learning (Montiel et al., 2020)  [^3] https://github.com/online-ml/river. We trained a support vector machine using the mean radius and mean texture and a Random Forest Classifier using mean concavity and mean symmetry. The mean radius of the features, the mean texture, the mean radius and the mean texture were chosen, so each model had features from one leg of the dendrogram into which the features of this data set fall into (Pantazi et al., 2002). These features are different from the features used in the non-streaming experiment. The support vector machine is $90\%$ accurate with $\alpha_{1}=FPR=0.18$ and $\beta_{2}=FNR=0.06$. The Random Forest is $88\%$ accurate with $\alpha_{2}=FPR=0.23$ and $\beta_{2}=FNR=0.09$.
@@
The categorizers then applied predictions to the online data and, for each time step, the contingency table was updated, the new sample was recorded, and latent profile analysis was applied to the new sample, plus the complete history of the samples. Because latent profiles are calculated when a new sample arrives, the latent class of a single sample will often flip. Due to this, Figure 6 shows the history of the parameters of interest calculated for the last latent profile. Additionally, due to the instability of the discriminant for Hui-Walter, the first $250$ time steps are omitted because they contain discontinuities. The literature on Hui-Walter states that Hui-Walter may be better at estimating the test agreement than the actual false positive and false negative rates (Bertrand et al., 2005), and the results of our experiment in Figure 6 show the parameters that come close to the desired metrics of the training environment scaled up or down by a factor of $2$ for $\alpha_{i},\beta_{i}$ for $i=1,2$. Prior probabilities $\theta_{1},\theta_{2}$ are more stable. Empirically, the priors are $\theta_{1}=0.11$ and $\theta_{2}=0.51$, calculated by taking the number of true values from the labels available in our experimental setup.
@@
Figure 6: Online-Hui Walter applied to the Wisconsin Breast Cancer data set in a stream with the river python library for online machine learning Montiel et al. (2020) to simulate online data and the balanced accuracy for both classifiers was calculated over two latent subpopulations starting after a burn in period of $250$ streaming samples.
@@
Table 9 contains the mean absolute error of the values estimated with Hui-Walter versus those estimated in the training environment. We calculate the mean absolute error using the last $200$ steps. This technique is similar to the burn-in period in Bayesian statistics.
@@
The Rand index (RI) is a measure of the similarity between two cluster assignments, typically the ground truth labels and the clustering results Rand (1971). It is calculated by considering the number of pairs of data points that are either in the same cluster in both assignments (true positives, TP) or in different clusters in both assignments (true negatives, TN). Given the total number of pairs of data points, N, the Rand index can be computed as follows.
@@
In clustering problems, the Rand index can be considered a measure of the "accuracy" of a clustering algorithm when compared to the ground truth labels. However, it is essential to note that there are some differences between these two measures.
@@
Table 7: Wisconsin Breast Cancer Data Set: Comparison of the mean Balanced Accuracy estimated by Hui-Walter on the online data, Rand Index, and Ground Truth Accuracy for Random Forest and Support Vector Machine
@@
In order to compare our method to established methodologies from unsupervised learning, we compare the balanced accuracy found for our online implementation of Hui-Walter with the Rand Index which has been shown to relate the accuracy to the unsupervised learning setting (Garg & Kalai, 2018). For the Wisconsin Breast Cancer data set, we compare the balanced accuracy for the support vector machine and the random forest to the Rand index for both classifiers over both latent sub-populations. For the random forest model, we found a Rand index of 0.78 and a balanced accuracy of 0.96 from Hui-Walter, and the ground truth accuracy is 0.86. For the support vector machine, the balanced accuracy is 0.73, the RI is 0.82, and the ground truth accuracy is 0.90. These can be seen in Table 7 where we can see that online Hui-Walter gives a clear improvement over perhaps the only available baseline metric for this problem. The results are even more compelling for the analogous experiment run on the adult data set; as we can see in Table 8, the Balanced Accuracy is closer to the ground truth accuracy than the Rand index. The classifiers in question (Random Forest, Support Vector Machine, and Logistic Regression) were all chosen for their simplicity and ubiquity in production settings.
Table 9: Mean Absolute Error for Streaming Parameters
@@
Although this test yielded reasonable estimates for most parameters, the main issue with this approach for online estimation is the need for more standard errors, which is given by Bayesian estimation for the static case (Johnson et al., 2001). The worst estimate on the streaming data, $\theta_{1}$, also closely matches the static data case found with the Gibbs sampling method.
Conclusion
Our results show that the Hui-Walter method works very well for static data and gives plausible results for online data, and this method requires further improvements. The data from this experiment are in a three-way contingency table, and a log-linear model could improve the estimate as is in (Hui & Zhou, 1998). Additionally, the algebraic geometry of the three-way contingency tables could yield different results, as there is a relationship between tensor factorization and the parameters of the product-multinomial model (Dunson & Xing, 2009). One of the limitations of using explicit formulas for the $2\times 2\times 2$ case is that sometimes the solutions do not have plausible solutions (Hui & Walter, 1980). In other words, it is possible to get a solution for the false positive rate greater than one or less than zero. Furthermore, the performance of the straightforward solutions on the online data lacked confidence intervals, which are problematic for applied settings, and the solutions are not as precise as the Gibbs sampler. One suggestion for further research on how this can be improved is to leverage online Gibbs sampling (Kim et al., 2016; Dupuy et al., 2017).
@@ Conclusion
Broader Impact Statement
This research is expected to significantly impact the field of applied machine learning by providing a methodology to assess the efficacy of the model in the common but challenging circumstances of incomplete, unknown, or dynamically changing data sets. The Hui-Walter method, despite its limitations, as highlighted in our study, offers a robust starting point for practitioners dealing with data ambiguity. Additionally, our explorations of potential improvements, such as the application of online Gibbs sampling, open a promising path towards refining this technique. Ultimately, the ability to gauge the effectiveness of machine learning models even in the absence of complete data will empower practitioners across industries, driving more confident and informed decision-making processes.

### v2

@@ 7.3 Summary of Experimental Results
Our experiments show that the model’s false positive and false negative rates and prior can be estimated on unknown data sets with the Hui-Walter method. Although we assumed the default distribution $Beta(1,1)$ for the errors and priors, the results from the two experiments are generally promising for rough order of magnitude estimates. With more iterations or a more specific distributional assumption, the results would improve. The results would also likely be more favorable with better trained models as well.
8 Hui-Walter Online
@@
Our main goal in this research is to answer how we estimate false positive and false negative rates of binary categorizers to a stream of unlabeled data without available ground truth. Current online machine learning theory still assumes that labels are available for the data in question [17]. Figure 6 shows an experiment where we streamed the Wisconsin Breast Cancer Data set with the river python library for online machine learning [27]  [^4] https://github.com/online-ml/river.
These features are different from the features used in the non-streaming experiment. The support vector machine is $90\%$ accurate with $\alpha_{1}=FPR=0.18$ and $\beta_{2}=FNR=0.06$. The Random Forest is $88\%$ accurate with $\alpha_{2}=FPR=0.23$ and $\beta_{2}=FNR=0.09$.
@@
The categorizers then applied predictions to the online data and, for each time step, the contingency table was updated, the new sample was recorded, and latent profile analysis was applied to the new sample, plus the complete history of the samples. Because latent profiles are calculated when a new sample arrives, the latent class of a single sample will often flip. Due to this, Figure 6 shows the history of the parameters of interest calculated for the last latent profile. Additionally, due to the instability of the discriminant for Hui-Walter, the first $250$ time steps are omitted because they contain discontinuities. The literature on Hui-Walter states that Hui-Walter may be better at estimating the test agreement than the actual false positive and false negative rates [4], and the results of our experiment in Figure 6 show the parameters that come close to the desired metrics of the training environment scaled up or down by a factor of $2$ for $\alpha_{i},\beta_{i}$ for $i=1,2$. Prior probabilities $\theta_{1},\theta_{2}$ are more stable. Empirically, the priors are $\theta_{1}=0.11$ and $\theta_{2}=0.51$, calculated by taking the number of true values from the labels available in our experimental setup.
@@
Figure 6: Online-Hui Walter applied to the Wisconsin Breast Cancer data set in a stream with the river python library for online machine learning [27] to simulate online data and the balanced accuracy for both classifiers was calculated over two latent subpopulations starting after a burn in period of $250$ streaming samples.
@@
Table 9 contains the mean absolute error of the values estimated with Hui-Walter versus those estimated in the training environment. We calculate the mean absolute error using the last $200$ steps. This technique is similar to the burn-in period in Bayesian statistics  [^5] Data and Experiments available: https://github.com/kslote1/hui-walter.
@@
The Rand index (RI) is a measure of the similarity between two cluster assignments, typically the ground truth labels and the clustering results [29]. It is calculated by considering the number of pairs of data points that are either in the same cluster in both assignments (true positives, TP) or in different clusters in both assignments (true negatives, TN). Given the total number of pairs of data points, N, the Rand index can be computed as follows.
@@
In clustering problems, the Rand index can be considered a measure of the “accuracy” of a clustering algorithm when compared to the ground truth labels. However, it is essential to note that there are some differences between these two measures.
@@
Table 7: Wisconsin Breast Cancer Data Set: Comparison of the mean Balanced Accuracy estimated by Hui-Walter on the online data, Rand Index, and Ground Truth Accuracy for Random Forest and Support Vector Machine.
@@
In order to compare our method to established methodologies from unsupervised learning, we compare the balanced accuracy found for our online implementation of Hui-Walter with the Rand Index which has been shown to relate the accuracy to the unsupervised learning setting [14]. For the Wisconsin Breast Cancer data set, we compare the balanced accuracy for the support vector machine and the random forest to the Rand index for both classifiers over both latent sub-populations. For the random forest model, we found a Rand index of 0.78 and a balanced accuracy of 0.96 from Hui-Walter, and the ground truth accuracy is 0.86. For the support vector machine, the balanced accuracy is 0.73, the RI is 0.82, and the ground truth accuracy is 0.90. These can be seen in Table 7 where we can see that online Hui-Walter gives a clear improvement over perhaps the only available baseline metric for this problem. The results are even more compelling for the analogous experiment run on the adult data set; as we can see in Table 8, the Balanced Accuracy is closer to the ground truth accuracy than the Rand index. The classifiers in question (Random Forest, Support Vector Machine, and Logistic Regression) were all chosen for their simplicity and ubiquity in production settings.
@@
Table 9: After streaming the data, the parameters that the model ended on compared to the known parameter values through Mean Absolute Error.
Although this test yielded reasonable estimates for most parameters, the main issue with this approach for online estimation is the need for more standard errors, which is given by Bayesian estimation for the static case [22]. The worst estimate on the streaming data, $\theta_{1}$, also closely matches the static data case found with the Gibbs sampling method.
9 Conclusions
Our results show that the Hui-Walter method works very well for static data and gives plausible results for online data, and this method requires further improvements. The data from this experiment are in a three-way contingency table, and a log-linear model could improve the estimate as is in [20]. Additionally, the algebraic geometry of the three-way contingency tables could yield different results, as there is a relationship between tensor factorization and the parameters of the product-multinomial model [10]. One of the limitations of using explicit formulas for the $2\times 2\times 2$ case is that sometimes the solutions do not have plausible solutions [19]. In other words, it is possible to get a solution for the false positive rate greater than one or less than zero. Furthermore, the performance of the straightforward solutions on the online data lacked confidence intervals, which are problematic for applied settings, and the solutions are not as precise as the Gibbs sampler. One suggestion for further research on how this can be improved is to leverage online Gibbs sampling [23, 11].
@@ 9 Conclusions
10 Limitations
The Hui-Walter method and the online methods mentioned in this text use data sets with even distributions of classes. Further work should be and considerations should be made on heavily unbalanced classes or classes where the prevalence of the phenomenon is very low.
For the Wisconsin Breast Cancer Data Set, it is essential to note that the data set has some limitations. For example, data are based on a specific population of patients and may not represent all cases of breast cancer. Furthermore, the data are from the 1990s and may not reflect more recent advances in breast cancer diagnosis and treatment. Despite these limitations, the Wisconsin breast cancer data set remains an essential resource for researchers studying breast cancer and working to improve the diagnosis and treatment of this disease. It has contributed significantly to our understanding of breast cancer and will continue to be a valuable resource for researchers in the field.
Acknowledgments
We would like to acknowledge Yichuan Zhao, Herbert Roitblatt, and Igor Belykh for their valuable feedback.